put the following equation into sturm-liouville form x^2y^''+xy^'+(λx^2-n^2)y=0
1
Expert's answer
2020-08-09T18:24:04-0400
"Solution"

Bessel’s equation

"x^2 y^{''}+xy^{'}+(\lambda x^2-n^2)y=0"


Theorem. Any second order linear operator can be put into the form of the Sturm-Liouville operator


This is in the correct form. We just identify "p(x) =a_2(x)" and "q(x) =a_0(x)" .However, considering the Bessel's equation;


"a_2(x) =x^2" and "a^′_2(x) = 2x \ne a_1(x)."


In the Sturm Liouville operator the derivative terms are gathered together into one perfect derivative


We need only multiply this equation by


"\frac{1}{x^2}\epsilon^{\int \frac{dx}{x}}=\frac{1}{x}"


to put the equation in Sturm-Liouville form;


"\frac{x^2 y^{''}}{x}+\frac{xy^{'}}{x}+(\frac{\lambda x^2}{x}-\frac{n^2}{x})y=x y^{''}+y^{'}+(\lambda x-\frac{n^2}{x}
)y=(xy^{'})^{'}+(\lambda x-\frac{n^2}{x})y=0"




"(xy^{'})^{'}+(\lambda x-n^2/x)y=0 ----->Answer"